# Spectral Portfolio Theory: From SGD Weight Matrices to Wealth Dynamics

### 09 Mar 2026 v1

@@ 3 Portfolio Dynamics: The Three Forces of SGD
$d\sigma_{k}=\Bigl[-\eta\,u_{k}^{\top}(\nabla_{W}\mathcal{L})\,v_{k}+\eta D\Bigl(\frac{m-n+1}{2\sigma_{k}}-\sum_{j\neq k}\frac{\sigma_{k}}{\sigma_{k}^{2}-\sigma_{j}^{2}}\Bigr)\Bigr]dt+\sqrt{2\eta D}\,d\beta_{k}.$ (3)
@@ 4 The Stationary Portfolio: Bulk and Tail
$p_{\sigma}(\sigma)\propto\sigma^{m-n+1}\,e^{-(\beta_{1}/4\eta D)\,\sigma^{2}},$ (7)
@@
  • Tail: a few large $\sigma_{k}$ values form a power-law tail ($\sigma^{m-n+1}$ for large $\sigma$). These are concentrated satellite positions—large bets on specific factors.
@@
The tail exponent mn+1m-n+1 depends on the aspect ratio of the allocation matrix. For a market with $N$ assets observed over $T$ periods, the exponent is TN+1T-N+1. More observation (larger $T$) fattens the tail (allows more concentration); more assets (larger $N$) thins it (forces diversification). This is testable against Norwegian portfolio data.
@@ 6.1 The Additive Regime: Short Horizon
This is also the regime where Olsen et al. (2025)’s spectral theory of SGD applies most directly: weight updates are small additive perturbations, the matrix-valued SDE is well approximated by Dyson Brownian motion, and the stationary spectral density follows a gamma-type distribution with power-law tails.
@@ 9 The Fokker–Planck Connection
The stationary solution has gamma-type density with power-law tail exponent mn+1m-n+1.
@@ 11.2 Spectral Consequences
    a scale-shifted version of the pre-perturbation distribution. The shape (the exponent mn+1m-n+1 in the power-law tail) is preserved.
@@ 13 The Spectral Invariance Conjecture
    The stationary singular-value distribution has tail exponent TN+1T-N+1, and the aggregate wealth distribution formed by $x=\|W\|_{F}$ has Pareto tail exponent $\alpha=1+v/D$ where $v$ is the mean drift of $\|W\|_{F}$.
@@
Interpretation. The Conjecture provides a unified criterion for evaluating any environmental perturbation—tax, regulation, transaction cost, or fee structure—in terms of its spectral signature. Part (a) connects spectral and Pareto exponents; part (b) characterises benign (isotropic) perturbations; part (c) quantifies distortion from differential treatment.
@@ 14.2 Score Matching and Denoising Score Matching
For a linear network $\hat{v}(x;W)=Wx$, the gradient is $\nabla_{W}\mathcal{L}=-\frac{2}{T}\sum_{t}r_{t}\,x_{t}^{\top}$, and the gradient covariance factors as $C=4\,\Sigma_{r}\otimes\Sigma_{x}/T$. This Kronecker structure is the input assumption of Olsen et al. (2025), Theorem 3.2 (spectral characterisation of SGD stationary distributions). The resulting spectral density of $W$ follows a gamma-type distribution with tail exponent mn+1m-n+1, which maps to a Pareto exponent $\alpha=1+f(m,n,\beta_{1}/\eta D)$ via (21).
@@ 16.1 General Spectral Predictions
    Spectral tail exponent from portfolio data. Compute the SVD of wealth portfolios (e.g. from Norwegian SSB microdata). The tail exponent of the singular-value distribution should be TN+1T-N+1 where $T$ is the observation window and $N$ the number of asset classes.

### 16 Apr 2026 v2

@@ 3 Portfolio Dynamics: The Three Forces of SGD
$d\sigma_{k}=\Bigl[-\eta\,u_{k}^{\top}(\nabla_{W}\mathcal{L})\,v_{k}+\eta D\Bigl(\frac{m-n+1}{2\sigma_{k}}+\sum_{j\neq k}\frac{\sigma_{k}}{\sigma_{k}^{2}-\sigma_{j}^{2}}\Bigr)\Bigr]dt+\sqrt{2\eta D}\,d\beta_{k}.$ (3)
@@ 4 The Stationary Portfolio: Bulk and Tail
$p_{\sigma}(\sigma)\propto\sigma^{(m-n+1)/2}\,e^{-(\beta_{1}/4\eta D)\,\sigma^{2}},$ (7)
@@
  • Tail: a few large $\sigma_{k}$ values form a power-law tail ($\sigma^{(m-n+1)/2}$ for large $\sigma$). These are concentrated satellite positions—large bets on specific factors.
@@
The tail exponent (mn+1)/2(m-n+1)/2 depends on the aspect ratio of the allocation matrix. For a market with $N$ assets observed over $T$ periods, the exponent is (TN+1)/2(T-N+1)/2. More observation (larger $T$) fattens the tail (allows more concentration); more assets (larger $N$) thins it (forces diversification). This is testable against Norwegian portfolio data.
@@ 6.1 The Additive Regime: Short Horizon
This is also the regime where Olsen et al. (2025)’s spectral theory of SGD applies most directly: weight updates are small additive perturbations, the matrix-valued SDE is well approximated by Dyson Brownian motion, and the stationary spectral density follows a gamma-type distribution with power-law tails. Aarts et al. (2025) provide independent confirmation from the physics side, showing that the eigenvalue dynamics of SGD weight matrices are described by Dyson Brownian motion with a Coulomb gas stationary distribution whose variance scales as $\alpha/|\mathcal{B}|$ (learning rate over batch size).
@@ 6.4 The Transition: Marchenko–Pastur to Inverse-Wishart
Remark (Generalised Dyson Brownian Motion).
Bousseyroux and Bouchaud (2025) extend the Dyson Brownian motion framework by interpreting the additivity of R-transforms in free probability as a dynamical evolution of eigenvalues interacting through two-body and higher-body forces. Their generalised Dyson equation accommodates motion matrices $\mathbf{B}$ with arbitrary free cumulants $\kappa_{n}^{\mathbf{B}}$—not just the Gaussian case ($n=2$) of standard DBM. When $\kappa_{n}^{\mathbf{B}}\neq 0$ for $n\geq 3$, the eigenvalue dynamics acquire $n$-body repulsion terms that modify both the bulk density and the conditions under which outlier eigenvalues detach. Since asset returns are heavy-tailed, this extension is relevant to the multiplicative DBM of equation (12): replacing Gaussian noise $\xi_{i}$ with $\alpha$-stable driving produces a different family of stationary spectral densities whose tail exponent depends on both the compounding horizon and the noise stability index $\alpha$. The free-convolution machinery also provides a natural algebraic setting for computing the spectral density of products of free random matrices—the operation that arises when portfolio allocations are composed across time periods or aggregated across investors.
@@ 9 The Fokker–Planck Connection
The stationary solution has gamma-type density with power-law tail exponent (mn+1)/2(m-n+1)/2.
@@ 10 The Bouchaud–Mézard Unification
The binary condensation/no-condensation dichotomy above admits a richer phase structure when agents have heterogeneous growth rates. Bernard et al. (2026) solve the mean-field version of the Bouchaud–Mézard model with quenched disorder in the drift $\eta_{i}$ and identify three phases as a function of the redistribution rate $\varphi$: (i) a localised phase ($\varphi<\varphi_{c}$), in which a single agent captures a macroscopic fraction of total wealth—the analogue of Bose–Einstein condensation; (ii) a partially localised phase, in which the wealth distribution has a power-law tail with exponent $\mu=1-\Sigma_{0}^{2}/\sigma^{4}$, where $\Sigma_{0}^{2}$ is the variance of the quenched growth rates and $\sigma^{2}$ the diffusion intensity; and (iii) a delocalised phase ($\varphi>\varphi_{c}$), in which wealth is spread across agents. The critical redistribution rate $\varphi_{c}$ marks a genuine phase transition, derived using an analogy with Derrida’s Random Energy Model. In the spectral language, the localised phase corresponds to a participation ratio $p_{1}\to 1-\varphi/\varphi_{c}$, so that a single eigenmode dominates the wealth vector—a spectral fingerprint of condensation that is absent in the delocalised phase.
@@ 11.2 Spectral Consequences
    a scale-shifted version of the pre-perturbation distribution. The shape (the exponent (mn+1)/2(m-n+1)/2 in the power-law tail) is preserved.
@@ 13 The Spectral Invariance Conjecture
    The stationary singular-value distribution has tail exponent (TN+1)/2(T-N+1)/2, and the aggregate wealth distribution formed by $x=\|W\|_{F}$ has Pareto tail exponent $\alpha=1+v/D$ where $v$ is the mean drift of $\|W\|_{F}$.
@@
Interpretation. The Conjecture provides a unified criterion for evaluating any environmental perturbation—tax, regulation, transaction cost, or fee structure—in terms of its spectral signature. Part (a) connects spectral and Pareto exponents; part (b) characterises benign (isotropic) perturbations; part (c) quantifies distortion from differential treatment. The free-convolution framework of Bousseyroux and Bouchaud (2025)—which interprets the additivity of R-transforms as a generalised Dyson evolution—provides a natural algebraic setting for attacking parts (a) and (b), since the spectral density of the allocation matrix at finite horizon is a free convolution of the initial condition with the noise contribution.
@@ 14.2 Score Matching and Denoising Score Matching
For a linear network $\hat{v}(x;W)=Wx$, the gradient is $\nabla_{W}\mathcal{L}=-\frac{2}{T}\sum_{t}r_{t}\,x_{t}^{\top}$, and the gradient covariance factors as $C=4\,\Sigma_{r}\otimes\Sigma_{x}/T$. This Kronecker structure is the input assumption of Olsen et al. (2025), Theorem 3.2 (spectral characterisation of SGD stationary distributions). The resulting spectral density of $W$ follows a gamma-type distribution with tail exponent (mn+1)/2(m-n+1)/2, which maps to a Pareto exponent $\alpha=1+f(m,n,\beta_{1}/\eta D)$ via (21).
@@ 16.1 General Spectral Predictions
    Spectral tail exponent from portfolio data. Compute the SVD of wealth portfolios (e.g. from Norwegian SSB microdata). The tail exponent of the singular-value distribution should be (TN+1)/2(T-N+1)/2 where $T$ is the observation window and $N$ the number of asset classes.